# NeRF-Supervised Feature Point Detection and Description

### v2

@@ 5.2 Relative Pose Estimation
We leverage the ScanNet dataset [15] to evaluate the relative pose estimation on indoor scenes. We employ the 1500 test image pairs provided by [23].
We employ the YFCC100M [16] and MegaDepth [17] datasets to assess pose estimation on outdoor scenes. We utilise the YFCC100M test split introduced by [23], this split samples 1000 image pairs from four scenes each, thus, the total number of image pairs for the outdoor pose evaluation is 4000 image pairs. Additionally, we use the MegaDepth1500 test split introduced by [21, 24]. This test split comprises 1500 image pairs from two phototourism scenes: “Sacre Coeur” and “St. Peter’s Square”.
@@
For indoor relative pose evaluation, we follow [24, 23] and resize ScanNet’s images to a size of [640 $\times$ 480]. For SuperPoint, we use the top 1k predicted points and 20k for SiLK.
Additionally, for outdoor relative pose evaluation, we resize the YFCC100M and MegaDepth1500 images so that the longest dimension is equal to 1200 and 1600 respectively. We use the top 2k predicted points for SuperPoint and 20k for SiLK.
For both indoor and outdoor pose evaluations, we report the pose error AUC at thresholds ($5^{\circ}$, $10^{\circ}$, $20^{\circ}$), where the pose error is the maximum between the angular rotation error and angular translation error. Furthermore, we compute the essential matrix using the matched interest points using OpenCV with a RANSAC threshold of 0.5 over the mean of the focal length.
@@
Similar to the Homography Estimation evaluation Sec. 5.1, we compare SuperPoint-PrP and SuperPoint-PrP-Hyb with the baseline SuperPoint model. Likewise, we assess SiLK-PrP and SiLK-PrP-Aug against the baseline SiLK model.
@@
As depicted in Tab.3, the PrP models consistently surpass their respective baseline counterparts across all angular pose error thresholds. Although performance enhancement for SuperPoint-PrP and SuperPoint-PrP-Hyb over SuperPoint [6] is marginal, a more substantial improvement is evident in the case of SiLK-PrP and SiLK-PrP-Aug in comparison to SiLK [14], particularly as the angular threshold increases.
Unfortunately, the relative pose estimation computes the translation error as the angular translation error between the ground truth and estimated translation vectors. The computed angular translation error suffers from instability as it could only be computed up to a certain scale factor [43, 44]. This instability could be seen in Fig. 4, where it is evident that the angular translation error is unstable when the norm of the ground truth relative translation vector ($||t_{GT}||$) is approximately 0.2 and below.
To address this issue, we refined the indoor relative pose estimation assessment. For scenes where $||t_{GT}||$ is below the threshold $\epsilon$ set at 0.15, we exclusively report the pose error AUC based on angular rotation error. Furthermore, for scenes where the threshold $\epsilon$ exceeds 0.15, we report the pose error AUC as the maximum value between angular rotation and translation errors. As presented in Tab. 4, it is evident that all models excel in estimating relative rotation in scenes characterised by minimal translation between viewpoints. Notably, the PrP models exhibit a slight performance advantage over their baseline counterparts in this specific scenario. Additionally, in scenes where the ground truth relative translation exceeds the defined threshold, further improvement can be observed for the PrP models compared to results reported in Tab. 3.
Figure 4: Instability in Angular Translation Error. Density plot illustrating how the angular translation error is unstable in situations where the ground truth relative translation between two camera viewpoints is minimal.
@@
Table 4: Separated Indoor Relative Pose Estimation Error AUC. For image pairs with minimal relative translation in viewpoints, the PrP models exhibit superior performance in accurately estimating the relative rotation, especially the SuperPoint-PrP and SiLK-PrP-Aug models.
@@
Similar to results obtained in Tab. 3, the PrP models surpass their corresponding baseline models in relative pose estimation for outdoor scenes, as observed by results shown in Tabs. 5 and 6. The results reported in Tabs. 3 to 6 collectively highlight that models subjected to PrP supervision during training demonstrate significant improvement in their relative pose estimation capability.
Table 5: YFCC Relative Pose Estimation
Error AUC.
@@
Table 6: MegaDepth Relative Pose Estimation
Error AUC.
@@ 5.3 Pairwise Point Cloud Registration
We follow recent work [45, 14] and evaluate the interest point detectors and descriptors on a 3D point cloud registration task. Evaluation is performed on ScanNet’s [15] official test split, where image pairs are sampled 20 frames apart.
@@
The point cloud registration evaluation is conducted by resizing the ScanNet [15] images to dimensions of [128 $\times$ 128]. Given a pair of RGB-D images, a 6-DOF pose is estimated that best aligns the first image to the second image. The metrics reported during the evaluation are angular rotation and translation errors in degrees and centimetres respectively. Moreover, the chamfer distance is also reported between the ground-truth point cloud and the reconstructed point cloud in centimetres. For all three reported metrics the accuracy, mean and median errors over the entire dataset are reported at different thresholds [45].
@@
Unlike previous evaluations (Secs. 5.1 and 5.2) where sparse feature matching was employed, we follow [45, 14] and perform the point cloud registration using dense feature matching. The ratio test is utilised to find correspondences and alignment is seen as a Procustes problem solved by a weighted Kabsch’s algorithm. [46, 47, 14].
@@ 6 Conclusion
This paper presented a novel approach to supervise feature point detectors and descriptors using perspective projective geometry on synthetic data. Despite using a smaller dataset, our PrP approach achieves similar or better performance without compromising generalisability. Moreover, we noticed that the PrP-supervised models converge faster and outperform homography-supervised models in non-planar multi-view benchmarks, though, they slightly lag in homography estimation. We emphasise the main advantage of the PrP approach is not only benchmark improvements, but also its ability to train on any custom multi-view synthetic dataset with less data and less training time. The larger potential for further advancements depends on enhancing neural rendering to produce higher-quality synthetic images and more accurate depth maps, reducing artefacts and misprojection errors.

### v3

@@ 5.2 Relative Pose Estimation
We leverage the ScanNet dataset [15] to evaluate the relative pose estimation for indoor scenes, employing the test split of 1500 image pairs provided by [23].
For outdoor relative pose estimation, the YFCC100M [16] and MegaDepth [17] datasets are utilised. We adopt the YFCC100M test split introduced by [23], which includes 1000 image pairs sampled from four scenes, totalling 4000 image pairs. Additionally, we use the MegaDepth1500 test split, introduced by [21, 24], which consists of 1500 image pairs from two phototourism scenes: “Sacre Coeur” and “St. Peter’s Square”.
@@
For indoor relative pose evaluation, we follow [24, 23], resizing ScanNet images to [640 $\times$ 480]. We use the top 1k predicted points for SuperPoint and 20k for SiLK.
For the outdoor pose evaluation, the YFCC100M and MegaDepth1500 images are resized so that their longest dimensions are 1200 and 1600 pixels, respectively. We use the top 2k points for SuperPoint and 20k for SiLK.
In both cases, we report the pose error AUC at thresholds ($5^{\circ}$, $10^{\circ}$, $20^{\circ}$), where pose error is the maximum between the angular rotation and translation errors. We also compute the essential matrix using OpenCV’s RANSAC with a threshold of 0.5 over the mean focal length.
@@
As in Sec. 5.1, we compare SuperPoint-PrP, SuperPoint-PrP-Hyb, SuperPoint-PrP-(2k) and SuperPoint-(10k COCO) with the baseline SuperPoint model, likewise, we assess SiLK-PrP and SiLK-PrP-Aug against the baseline SiLK model.
@@
As depicted in Tab. 3, the PrP models consistently outperform their baselines across all angular pose error thresholds. While SuperPoint-PrP and SuperPoint-PrP-Hyb show only slight improvement over SuperPoint [6], SiLK-PrP and SiLK-PrP-Aug display a more notable gain, especially at higher angular thresholds. Additionally, SuperPoint-PrP-(2k) and SuperPoint-(10k COCO) follow the same trend as in Sec. 5.1, where they underperform compared to both baseline SuperPoint and SuperPoint-PrP, with SuperPoint-(10k COCO) displaying the weakest performance.
Unfortunately, the relative pose estimation computes the translation error as the angular translation between ground truth and estimated translation vectors. However, this metric is unstable due to scale factor ambiguity [43, 44]. As shown in Fig. 4, the angular translation error becomes unstable when the norm of the ground truth relative translation vector ($||t_{GT}||$) is approximately 0.2 or less.
To address this issue, we refined the indoor relative pose estimation assessment. For scenes where $||t_{GT}||$ is below 0.15, we only report pose error AUC based on angular rotation error. For scenes with $||t_{GT}||$ above 0.15, we use the maximum of angular rotation and translation errors. As presented in Tab. 4, all models excel at estimating the angular rotation in scenes with minimal translation between viewpoints, with the PrP models slightly outperforming their baselines. In scenes where $||t_{GT}||$ exceeds the defined threshold, the PrP models show further improvement compared to results reported in Tab. 3, while SuperPoint-PrP-(2k) and SuperPoint-(10k COCO) continue to lag, with the latter performing the worst.
Figure 4: Angular Translation Error Instability. Density plot illustrating how the angular translation error is unstable in situations when the ground truth relative translation between two camera viewpoints is minimal.
@@
Table 4: Separated Indoor Relative Pose Estimation Error AUC. For image pairs with minimal relative translation in viewpoints, the PrP models exhibit superior performance in accurately estimating the relative rotation, especially SuperPoint-PrP and SiLK-PrP-Aug.
@@
As seen in Tabs. 5 and 6, the PrP models outperform their baseline counterparts in outdoor relative pose estimation, mirroring the indoor results. While, SuperPoint-PrP-(2k) and SuperPoint-(10k COCO) continue to underperform compared to other models. The results reported in Tabs. 3 to 6 consistently highlight that the PrP-supervised models show significant improvement in relative pose estimation capabilities.
Table 5: YFCC Relative Pose
Estimation Error AUC.
@@
Table 6: MegaDepth Relative Pose
Estimation Error AUC.
@@ 5.3 Pairwise Point Cloud Registration
We follow recent work [45, 14] and evaluate the interest point detectors and descriptors on a 3D point cloud registration task using ScanNet’s [15] official test split. For the evaluation, image pairs are sampled 20 frames apart [45].
@@
For point cloud registration evaluation, ScanNet [15] images are resized to [128 $\times$ 128]. A 6-DOF pose is estimated to align pairs of RGB-D images. We report angular rotation error (in degrees), translation error (in centimetres), and chamfer distance between ground truth and reconstructed point clouds (in centimetres). The mean and median errors are reported across the entire dataset, as well as the accuracy at different thresholds [45].
@@
Unlike previous evaluations (Secs. 5.1 and 5.2) where sparse feature matching was employed, we follow [45, 14] and use dense feature matching for the point cloud registration evaluation. The ratio test is utilised to find correspondences, while alignment is seen as a Procustes problem solved by a weighted Kabsch’s algorithm. [46, 47, 14].
@@ 6 Conclusion
This paper presented a novel approach to supervise feature point detectors and descriptors using perspective projective geometry on synthetic data. Despite using a smaller dataset, our PrP approach achieves similar or better performance without compromising generalisability. Moreover, we noticed that the PrP-supervised models converge faster and outperform homography-supervised models in non-planar multi-view benchmarks, though, they slightly lag in homography estimation. We emphasise the main advantage of the PrP approach is not only benchmark improvements, but also its ability to train on any custom multi-view synthetic dataset with less data and training time. Additionally, while the real images used to render NeRF scenes could potentially train the detectors directly with COLMAP-produced data, this approach is limited. The NeRF dataset uses an average of 80-100 images per scene, totalling less than 1000 images with limited viewpoint diversity. The degradation in performance observed from SuperPoint-PrP-(2K) compared to SuperPoint-PrP suggests that training on even fewer real images would further impair model performance. The larger potential for further advancements depends on enhancing neural rendering to produce higher-quality synthetic images and more accurate depth maps, reducing artefacts and misprojection errors.
Acknowledgements
This work was partly conducted during an MSc in Robotics and Computation at the Department of Computer Science, University College London (UCL).